# BECS: A Privacy-Preserving Computing Sharing Mechanism for 6G Computing Power Network

### v2

@@ Abstract
The evolution towards AI-based 6G networks, building upon the connectivity foundation of 5G, promises unprecedented intelligent applications. However, realizing this vision necessitates a significant evolution in computing to effectively leverage distributed computing resources. This demand drives the development of the 6G Computing Power Network (CPN), which aims to aggregate and orchestrate heterogeneous computing resources spanning user devices, edge nodes, and cloud infrastructures. In this article, we propose BECS, a computing sharing mechanism based on evolutionary algorithm and blockchain, designed to balance task offloading among computing resources within 6G CPN, thereby enhancing the computing resource utilization. We model computing sharing as a multi-objective optimization problem, aiming to improve resource utilization while addressing other trade-offs. To tackle this NP-hard problem, we devise a kernel distance-based dominance relation and incorporate it into the Non-dominated Sorting Genetic Algorithm III, significantly enhancing population diversity. In addition, we propose a pseudonym scheme based on zero-knowledge proofs to protect user privacy during computing sharing. Finally, security analysis and simulation results demonstrate that BECS can effectively utilize all computing resources in the 6G CPN, significantly improving resource utilization while preserving user privacy.
Index Terms:
Computing sharing, 6G networks, computing power network (CPN), evolutionary algorithm, blockchain, pseudonym scheme.
@@ I Introduction
In the upcoming 6G era, it is anticipated that everything will be intelligently connected, enabling a wide range of data-intensive applications. By deeply integrating communication networks with various vertical industries, 6G will enable unprecedented applications such as Holographic Integrated Sensing and Communication (HISC), Artificial General Intelligence (AGI), and Digital Twins (DT) [1]. This indicates that AI will be one of the most crucial technologies for constructing a comprehensively intelligent 6G networks, enabling network services to evolve dynamically and autonomously in response to demand. The coordination of end-edge-cloud computing devices to form a Computing Power Network (CPN) is expected to become a leading paradigm for supporting ubiquitous intelligent services in 6G networks [2]. In essence, the CPN envisions intelligent orchestration, where computing tasks are dynamically offloaded to the most suitable execution environment based on real-time service demands. This fosters a highly dynamic and hybrid computing environment that promotes the sharing of computing resources across 6G networks.
@@
While Multi-access Edge Computing (MEC) in 5G networks paved the way for offloading computing tasks to the network edge and unleashing potential for the Internet of Things (IoT) [3], its limitations become apparent when facing the demands of the 6G era. MEC is typically characterized by localized deployments, limited resources, and siloed management, all of which hinder the integration and orchestration of ubiquitously distributed, heterogeneous computing resources across the network. In the 6G era, networks are expected to scale up to support trillions of connections. Meanwhile, intelligent and diversified applications will not only require computing resources that far exceed previous levels but will also urgently demand a ubiquitous computing paradigm. This paradigm should enable real-time coordination of network-wide computing resources, dynamic on-demand orchestration, and seamless cross-domain resource allocation [4]. As a result, 5G MEC is inadequate for meeting future computing demands in terms of both scale and operational paradigms. Therefore, 6G networks require the CPN to deeply integrate and intelligently orchestrate dispersed, heterogeneous, and geographically distributed computing resources across the entire network. Such integration promotes full resource utilization and enables efficient, large-scale computing sharing [5].
@@
Driven by the unprecedented features and services of future 6G networks, communication networks are increasingly exploring distributed or multi-center management models [6]. Furthermore, the 6G CPN aims to comprehensively mobilize vast heterogeneous devices, edge nodes, and cloud computing resources across organizational boundaries. This cross-domain and heterogeneous environment, which lacks a central authority of trust, poses significant challenges for the access, scheduling, trading, and management of computing resources [7]. Traditional centralized computing resource management architectures struggle to adapt to such large-scale, dynamically evolving distributed systems. These systems are not only inefficient but also vulnerable to single points of failure. Therefore, it is necessary to explore novel distributed network management and governance mechanisms to ensure that entities within the CPN interact and collaborate in an autonomous, reliable, and efficient manner. Blockchain, with its inherent decentralization and its ability to enhance trust and transparency among multiple stakeholders, emerges as a key enabling technology for future 6G network management [8]. It offers secure, transparent, and traceable solutions for computing sharing, trusted transactions, access control, and other critical functionalities within the CPN.
Furthermore, the ubiquitous sharing and collaboration of computing resources exacerbate the risk of user privacy leakage. The 6G CPN involves a wide variety of computing devices. In particular, when user devices participate in computing sharing, information such as user identity, behavioral patterns, and interaction records may be collected and analyzed by attackers to infer sensitive data or perform identity tracking [9]. These potential threats may discourage users from actively engaging in computing sharing. Therefore, to achieve efficient computing sharing, it is essential to implement secure and effective privacy protection strategies to safeguard user rights and establish a trustworthy CPN environment.
@@
To address these intertwined challenges of efficiency, trust, and privacy, this article introduces a computing allocation and trading mechanism for 6G CPN, named BECS, which utilizes Blockchain and Evolutionary algorithms for efficient Computing resource Sharing. Furthermore, it provides a pseudonym scheme to protect users’ privacy. Specifically, the main contributions of this article are summarized as follows:
@@
  • We propose a dynamic and efficient blockchain-based mechanism for computing sharing, aimed at ensuring secure allocation and trading of resources between any devices in 6G CPN, thus enhancing utilization.
@@
  • We formulate computing allocation as a multi-objective optimization problem (MOOP) with six objectives, employing an evolutionary algorithm to balance the interplay among these objectives, thus achieving optimal allocation schemes for 6G CPN.
@@
  • We design a novel pseudonym scheme based on the Schnorr protocol, which protects user privacy during computing sharing in 6G CPN.
The remainder of this article is organized as follows. Section II reviews related work. Section III introduces the system model and formulates the computing sharing problem. Section IV describes NSGA-III-KDR and its application in solving the computing allocation MOOP. Section V presents the proposed pseudonym scheme and computing trading. Section VI analyzes the security and computational complexity of the proposed scheme. Section VII presents simulation results. Section VIII concludes the article.
@@ II Related Works
In this section, we present computing allocation and sharing in CPN/end-edge-cloud, and elaborate blockchain for decentralized management and privacy preservation.
II-A Computing Allocation and Sharing in CPN/End-Edge-Cloud
@@
Therefore, formulating computing allocation as a MOOP has become a more effective approach for handling these conflicting objectives. Meanwhile, evolutionary algorithms are widely adopted due to their ability to identify a set of non-dominated solutions representing different trade-offs. Peng et al. [11] formulated complex task offloading in the IIoT as a four-objective MOOP and developed a method to dynamically allocate computing resources based on the NSGA-III. Gong et al. [12] employed the multi-objective evolutionary algorithm based on decomposition (MOEA/D) to optimize a three-objective edge task offloading problem, aiming to minimize delay and maximize rewards. While these approaches demonstrate the applicability of MOEAs, standard algorithms like NSGA-III and MOEA/D can face challenges when dealing with many-objective optimization, often struggling to maintain sufficient population diversity alongside convergence pressure [13]. This limitation can hinder the exploration of the full solution space, particularly for complex 6G CPN scenarios involving numerous performance dimensions and heterogeneous resources.
II-B Blockchain for Decentralized Management and Privacy Preservation
@@
Building on the need for decentralized trust and management in 6G networks, blockchain technology has been actively explored as a key enabler [14]. Xie et al. [15] exploited the immutability of blockchain to propose a resource trading mechanism based on sharding and directed acyclic graphs for large-scale 6G networks, thereby enhancing resource utilization efficiency. Nguyen et al. [16] proposed a blockchain-based mobile edge-cloud computation offloading scheme, leveraging the distributed characteristics of blockchain to to provide secure and trusted computing services. Wang et al. [17] proposed a provable secure blockchain-based federated learning framework for wireless CPN, aimed at accelerating the convergence of federated learning and enhancing the efficiency of wireless CPN. These works highlight blockchain’s potential to automate, secure, and streamline interactions like resource discovery, access control, scheduling coordination, and payment settlement via smart contracts [18], thereby fostering a reliable environment for large-scale computing sharing.
@@
6G networks will integrate AI to fully merge the physical and digital worlds, necessitating enhanced security and privacy [19]. Nguyen et al. [20] utilized blockchain to provide adequate security for task offloading in mobile edge computing. Wang et al. [21] proposed a blockchain-enabled cyber-physical system integrating cloud and edge computing to achieve secure computing offloading. Samy et al. [22] introduced a blockchain-based framework for task offloading, ensuring security, integrity, and privacy in mobile edge computing. Although leveraging the characteristics of blockchain can provide preliminary security and privacy protection for edge and cloud computing allocation and trading, device-to-device computing sharing in 6G networks will require more comprehensive solutions to ensure the security and privacy between devices [23].
@@
Distinct from the aforementioned works, BECS is designed as a holistic mechanism for the 6G CPN. It uniquely combines a blockchain architecture, advanced multi-objective optimization techniques, and cryptographic privacy protection to tackle the intertwined challenges of efficiency and security in computing sharing of 6G CPN.
@@ III System Model and Problem Formulation
In this section, we introduce the computing sharing system model considered by BECS, along with other models used in constructing the MOOP, including the communication, computing, and service models.
@@ III-A System Model
BECS is a generalized architecture that supports various types of computing sharing in 6G CPN, as illustrated in Fig. 1. All computing devices, labeled as $\mathbb{D}=\{D_{1},\ldots,D_{d},\ldots,D_{D}\}$, are classified into three layers: user computing layer, edge computing layer, and cloud computing layer. Specifically, the set of user computing devices, denoted as $\mathbb{U}=\{U_{1},\ldots,U_{u},\ldots,U_{U}\}$, includes smartphones, computers, wearable devices, IoT devices, vehicles, and other devices that directly interact with users. These devices typically possess limited computing power. The set of edge computing devices, denoted as $\mathbb{E}=\{E_{1},\ldots,E_{e},\ldots,E_{E}\}$, includes edge servers, roadside units, cloudlets, and other devices capable of providing time-sensitive computing services to users. The set of cloud computing devices, denoted as =$\mathbb{C}=\{C_{1},\ldots,C_{c},\ldots,C_{C}\}$, consists of remote computing centers capable of providing large-scale computing power. The evolutionary algorithm matches computing resource requesters with providers, facilitating a multi-dimensional measurement of the deep reuse of computing resources. Additionally, with the support of permissioned blockchain, BECS enables dynamic management and trading of fine-grained computing resources. The proposed architecture consists of three main components:
@@
This study primarily focuses on efficient computing sharing. Accordingly, the computing device $D_{d}$ is denoted as $D_{d}=\{\varsigma_{d},\psi_{d}\}$, where $\varsigma_{d}$ is usually measured by device’s CPU [24]. Meanwhile, $\mathbb{T}=\{T_{1},\ldots,T_{t},\ldots,T_{T}\}$ denotes the set of all computing tasks. The tuple $T_{t}=\{\varphi_{t},\xi_{t},\tau_{t},d_{t}\}$ describes the computing task $T_{t}$. $\varphi_{t}$ and $\xi_{t}$ can be obtained using methods described in [25], such as graph analysis. Notations that will be used are presented in Table I.
TABLE I: Summary of Notations
@@
$\alpha_{d}$ the tasks processed number per unit of $D_{d}$’s computing capacity

### v3

# BECS: A Privacy-Preserving Computing Resource Sharing Mechanism for 6G Computing Power Network

@@ Abstract
The 6G Computing Power Network (CPN) is envisioned to orchestrate vast, distributed computing resources for future intelligent applications. However, achieving efficient, trusted, and privacy-preserving computing resource sharing in this decentralized environment poses significant challenges. To address these intertwined issues, this article proposes a holistic blockchain and evolutionary algorithm-based computing resource sharing (BECS) mechanism. BECS is designed to dynamically and adaptively balance task offloading among computing resources within the 6G CPN, thereby enhancing resource utilization. We model computing resource sharing as a multi-objective optimization problem, aiming to navigate these trade-offs. To tackle this NP-hard problem, we devise a kernel-distance-based dominance relation and incorporate it into the Non-dominated Sorting Genetic Algorithm III (NSGA-III), thereby significantly enhancing population diversity. In addition, we propose a pseudonym scheme based on zero-knowledge proofs to protect user privacy during computing resource sharing. Finally, security analysis and simulation results demonstrate that BECS can effectively leverage all computing resources in the 6G CPN, thereby significantly improving resource utilization while preserving user privacy.
@@ I Introduction
In the upcoming 6G era, it is anticipated that everything will be intelligently connected, enabling a wide range of data-intensive applications. By deeply integrating communication networks with various vertical industries, 6G will enable unprecedented applications such as Holographic Integrated Sensing and Communication (HISC), Artificial General Intelligence (AGI), and Digital Twins (DT) [1, 65, 66]. This indicates that AI will be one of the most crucial technologies for constructing a comprehensively intelligent 6G networks, enabling network services to evolve dynamically and autonomously in response to demand. The coordination of end-edge-cloud computing devices to form a Computing Power Network (CPN) is expected to become a leading paradigm for supporting ubiquitous intelligent services in 6G networks [2]. In essence, the CPN envisions intelligent orchestration, where computing tasks are dynamically offloaded to the most suitable execution environment based on real-time service demands. This fosters a highly dynamic and hybrid computing environment that promotes the sharing of computing resources across 6G networks. However, realizing this vision in the 6G CPN presents fundamental challenges. In particular, it is necessary to design a resource sharing mechanism that can simultaneously ensure trustworthiness and preserve user privacy, while navigating the complex performance trade-offs inherent in such a massive and decentralized system.
@@
While Multi-access Edge Computing (MEC) and fog computing in 5G networks paved the way for offloading computing tasks to the network edge and unleashing potential for the Internet of Things (IoT) [47], their limitations become evident when faced with more complex demands in the future. These paradigms are typically characterized by localized deployments, hierarchical interactions, and siloed management, all of which hinder the seamless integration and orchestration of ubiquitously distributed, heterogeneous computing resources across the network. In the 6G era, networks are expected to scale up to support trillions of connections. Meanwhile, intelligent and diversified applications will not only require computing resources that far exceed previous levels but will also urgently demand a ubiquitous computing paradigm. This paradigm must enable real-time coordination of network-wide computing resources and dynamic on-demand orchestration. Crucially, it involves navigating complex trade-offs among multiple, often conflicting, performance objectives such as minimizing latency, reducing energy consumption, and maximizing resource utilization across seamless cross-domain allocations [48]. The high-dimensional and strongly-coupled nature of these objectives presents a formidable optimization challenge, pushing standard multi-objective algorithms towards their performance limits, particularly in maintaining the solution diversity required to map the entire Pareto-optimal front. As a result, 5G MEC and fog computing are inadequate for meeting future computing demands in terms of both scale and operational paradigms. Therefore, 6G networks require the CPN to deeply integrate and intelligently orchestrate dispersed, heterogeneous, and geographically distributed computing resources across the entire network. Such integration promotes full resource utilization and enables efficient, large-scale computing resource sharing through collaborative offloading and incentive mechanisms [67, 68, 69].
@@
Driven by the unprecedented features and services of future 6G networks, communication networks are increasingly exploring distributed or multi-center management models [49]. Furthermore, the 6G CPN aims to comprehensively mobilize vast heterogeneous devices, edge nodes, and cloud computing resources across organizational boundaries. This cross-domain and heterogeneous environment, which lacks a central authority of trust, poses significant challenges for access, scheduling, trading, and management of computing resources [50]. Traditional centralized architectures not only struggle with inefficiency and single points of failure, but they also lack the transparency and immutability needed to build trust among diverse stakeholders. Therefore, it is necessary to establish a novel distributed management mechanism that enables autonomous, reliable, and efficient collaboration among devices. Such a mechanism must provide verifiable, tamper-proof records for resource trading and enforce operational rules transparently, without relying on a central intermediary. Blockchain, with its inherent decentralization and its ability to enhance trust and transparency among multiple stakeholders, emerges as a key enabling technology for future 6G network management [51, 70, 71]. It offers secure, transparent, and traceable solutions for computing resource sharing, trusted trading, access control, and other critical functionalities within the CPN.
Furthermore, the ubiquitous sharing and collaboration of computing resources exacerbate the risk of user privacy leakage. Distinct from the localized hierarchical structure of fog computing, the 6G CPN orchestrates a wide variety of computing devices into a unified, network-wide computing continuum. In particular, when user devices participate in computing resource sharing, not only their identity but also their behavioral patterns and interaction records can be collected and linked, allowing attackers to infer sensitive information [52, 86, 87]. These potential threats create a significant barrier to participation. To achieve efficient computing resource sharing, it is essential to provide a mechanism that allows users to prove their legitimacy and conduct transactions without revealing their persistent identities, thus ensuring both accountability and anonymity.
@@
To address these intertwined challenges of efficiency, trust, and privacy, this article introduces a novel mechanism for 6G CPN, named BECS, which utilizes Blockchain and Evolutionary algorithms for efficient Computing Sharing. It is designed to create a robust and efficient ecosystem where heterogeneous computing resources can be shared on a large scale. By effectively tackling the critical issues of resource allocation efficiency, decentralized trust, and user privacy preservation, BECS aims to unlock the full potential of distributed computing in 6G networks. Performance evaluations confirm that the proposed BECS mechanism improves average resource utilization by 56.94% compared to an NSGA-III-based mechanism. Specifically, the main contributions of this article are summarized as follows:
@@
  • We propose a dynamic and efficient blockchain-based mechanism for computing resource sharing, aimed at ensuring secure allocation and trading of resources between any devices in 6G CPN, thus enhancing utilization.
@@
  • We formulate computing allocation as a comprehensive multi-objective optimization problem (MOOP) with six objectives, employing an evolutionary algorithm to balance the interplay among these objectives, thus achieving optimal allocation schemes for 6G CPN.
@@
  • We design a novel pseudonym scheme based on the Schnorr protocol, which protects user privacy during computing resource sharing in 6G CPN.
The remainder of this article is organized as follows. Section II reviews related work. Section III introduces the system model and formulates the computing resource sharing problem. Section IV describes NSGA-III-KDR and its application in solving the computing allocation MOOP. Section V presents the proposed pseudonym scheme and computing trading. Section VI analyzes the security and computational complexity of the proposed scheme. Section VII presents simulation results. Section VIII concludes the article.
@@ II Related Works
In this section, we first introduce the paradigm of the 6G CPN and contrast it with 5G MEC. Then, we discuss existing works on computing allocation and sharing. Finally, we elaborate blockchain for decentralized management and privacy preservation.
II-A The Paradigm of 6G Computing Power Network
The trajectory from 5G to 6G signals a profound transformation, moving beyond the traditional pursuit of enhanced communication metrics. Driven by the rapid advancement of AI, it represents a fundamental paradigm shift toward the deep integration of computing and networking [72, 73, 74]. Although 5G MEC optimizes computing services by pushing computing power to the network edge, its inherent limitations in scale and cross-domain orchestration prevent it from meeting the ubiquitous intelligence demands of 6G [55].
6G CPN will deeply integrate all computing devices within the network to form a distributed and intelligent computing environment [12]. By seamlessly abstracting, virtualizing, and integrating decentralized computing resources managed by different operators and originating from diverse device types, 6G CPN aims to unify the heterogeneous resource landscape. It constructs a unified computing architecture that spans from user devices to multi-domain edge nodes and cloud centers, capable of supporting complex service architectures [57]. This transition from isolated resources to a unified architecture also assigns different identity roles to computing devices within the network. In 5G MEC or fog computing, devices often operate within fixed hierarchies where user devices are typically resource requesters, while edge/fog nodes act as resource providers. In contrast, 6G CPN breaks this limitation and enables a flat, collaborative ecosystem where every computing device can function both as a requester and a provider, thereby enabling full utilization of all computing resources in the network [58].
II-B Computing Allocation/Sharing in CPN/End-Edge-Cloud
@@
TABLE I: Differences between BECS and other main related works
Ref. Core Scenario Optimization Method MOOP Cloud-Edge-End Collaboration Blockchain Features Advanced Privacy Scheme 6G Relevance
[2] Energy-efficient task transfer in Wireless CPN Multi-Agent Deep Reinforcement Learning × × × ×
[3] Efficient task offloading in Edge CPN Two-Stage Evolutionary Search × × × ×
[6] Computation offloading in Industrial IoT NSGA-III × × ×
[7] Dependent task offloading in MEC MOEA/D × × ×
[53] Secure computation offloading in IoT Deep Reinforcement Learning × × ×
[14] Cooperative task offloading in MEC Multi-Agent DRL, Game Theory × × × ×
[15] Secure computation offloading in cyber-physical systems Deep Reinforcement Learning × × ×
[16] Secure task offloading in MEC Distributed Deep Q-Learning × × ×
BECS Privacy-preserving computing resource sharing in 6G CPN NSGA-III-KDR
@@
Therefore, formulating computing allocation as a MOOP has become a more effective approach for handling these conflicting objectives. Meanwhile, evolutionary algorithms are widely adopted due to their ability to identify a set of non-dominated solutions representing different trade-offs. Peng et al. [6] formulated complex task offloading in the IIoT as a four-objective MOOP and developed a method to dynamically allocate computing resources based on the NSGA-III. Gong et al. [7] employed the multi-objective evolutionary algorithm based on decomposition (MOEA/D) to optimize a three-objective edge task offloading problem, aiming to minimize delay and maximize rewards. While these approaches demonstrate the applicability of MOEAs, standard algorithms like NSGA-III and MOEA/D face significant challenges when dealing with many-objective optimization, often struggling to maintain sufficient population diversity alongside convergence pressure [27]. This limitation becomes particularly pronounced in the context of our 6G CPN model, where the six-objective formulation and inherent resource heterogeneity necessitate a considerably richer and more diverse set of trade-off solutions than those typically yielded by standard algorithms. Consequently, although these existing tools are capable of generating feasible solutions, their ability to explore the full spectrum of high-quality and well-distributed trade-off strategies remains limited in such a demanding environment.
II-C Blockchain for Decentralized Management and Privacy Preservation
@@
Building on the need for decentralized trust and management in 6G networks, blockchain technology has been actively explored as a key enabler [9, 75, 80]. Xie et al. [10] exploited the immutability of blockchain to propose a resource trading mechanism based on sharding and directed acyclic graphs for large-scale 6G networks, thereby enhancing resource utilization efficiency. Nguyen et al. [53] proposed a blockchain-based mobile edge-cloud computation offloading scheme, leveraging the distributed characteristics of blockchain to provide secure and trusted computing services. Wang et al. [12] proposed a provable secure blockchain-based federated learning framework for wireless CPN, aimed at accelerating the convergence of federated learning and enhancing the efficiency of wireless CPN. These works highlight blockchain’s potential to automate, secure, and streamline interactions like resource discovery, access control, scheduling coordination, and payment settlement via smart contracts [54], thereby fostering a reliable environment for large-scale computing resource sharing.
@@
6G networks will integrate AI to fully merge the physical and digital worlds, necessitating enhanced security and privacy for robust human-to-virtual connectivity [13, 76, 77]. Nguyen et al. [14] utilized blockchain to provide adequate security for task offloading in mobile edge computing. Wang et al. [15] proposed a blockchain-enabled cyber-physical system integrating cloud and edge computing to achieve secure computing offloading. Samy et al. [16] introduced a blockchain-based framework for task offloading, ensuring security, integrity, and privacy in mobile edge computing. Although leveraging the characteristics of blockchain can provide preliminary security and privacy protection for edge and cloud computing allocation and trading, device-to-device computing resource sharing in 6G networks will require more comprehensive solutions to ensure the security and privacy between devices [17].
@@
An overview of related works is given in Table I. Distinct from the aforementioned works, BECS is designed as a holistic mechanism for the 6G CPN. It uniquely combines a blockchain architecture, advanced multi-objective optimization techniques, and cryptographic privacy protection to tackle the intertwined challenges of efficiency and security in computing resource sharing of 6G CPN.
@@ III System Model and Problem Formulation
In this section, we introduce the computing resource sharing system model considered by BECS, along with other models used in constructing the MOOP, including the communication, computing, and service models.
@@ III-A System Model
To meet the demands of future intelligent applications, BECS introduces a generalized architecture, as illustrated in Fig. 1, designed to support diverse forms of computing resource sharing in 6G CPN. In this architecture, all computing devices, denoted as $\mathbb{D}=\{D_{1},\ldots,D_{d},\ldots,D_{D}\}$, are categorized into three layers: the user computing layer, edge computing layer, and cloud computing layer. This classification fundamentally redefines their interplay. Moreover, BECS overcomes the rigid and siloed structure of traditional 5G MEC. Instead of treating computing resources as isolated within discrete physical layers, BECS abstracts and virtualizes all participating devices across the network into an integrated, unified, and orchestrated resource pool. Consequently, the layer classification reflects the functional capabilities of devices rather than enforcing rigid operational boundaries. The resources in $\mathbb{D}$ are inherently heterogeneous, encompassing devices from different administrative domains, multiple operators, and diverse ownerships, thereby forming a cohesive, network-wide computing continuum that can be dynamically orchestrated.
The proposed BECS is primarily designed for a metropolitan area network scenario. In this context, numerous computing devices are geographically distributed across the city and interconnected through a high-bandwidth, low-latency optical fiber backbone. This dense and high-speed infrastructure constitutes a key feature of future 6G-enabled smart city deployments [60]. Crucially, this architecture implies that the main communication bottlenecks and the most significant latency variations primarily arise from wireless access links and task processing, rather than from the highly stable, low-latency wired backbone. Accordingly, the BECS model is strategically designed to optimize these dominant factors. Specifically, the set of user computing devices, denoted as $\mathbb{U}=\{U_{1},\ldots,U_{u},\ldots,U_{U}\}$, includes smartphones, computers, wearable devices, IoT devices, vehicles, and other devices that directly interact with users. These devices typically possess limited computing power. The set of edge computing devices, denoted as $\mathbb{E}=\{E_{1},\ldots,E_{e},\ldots,E_{E}\}$, includes edge servers, roadside units, cloudlets, and other devices capable of providing time-sensitive computing services to users. The set of cloud computing devices, denoted as $\mathbb{C}=\{C_{1},\ldots,C_{c},\ldots,C_{C}\}$, consists of remote computing centers capable of providing large-scale computing power. The evolutionary algorithm matches computing resource requesters with providers, facilitating a multi-dimensional measurement of the deep reuse of computing resources. Additionally, with the support of permissioned blockchain, BECS enables dynamic management and trading of fine-grained computing resources. The proposed architecture consists of three main components:
@@
In contrast to the static client-server architecture of 5G MEC, CRP and CRR in BECS do not represent fixed device roles but are instead transient functional states of computing devices, determined by their real-time computational surplus or deficit. This fluidity and flexibility are essential to unlocking the potential of computing resource sharing in 6G CPN. Meanwhile, with the support of blockchain and evolutionary algorithms, computing devices can autonomously match and schedule supply and demand, thereby meeting the ubiquitous and heterogeneous computing demands anticipated in future networks. To ensure this sharing is always productive and aligns with the network’s goal of enhancing overall computational capability, the BECS mechanism is founded on the principle that tasks are offloaded to devices with superior capacity. This approach prevents non-productive resource cycling and stabilizes the market, thereby ensuring that sharing is driven by genuine computing resource need and ultimately fostering a trusted ecosystem.
This study primarily focuses on efficient computing resource sharing. Accordingly, the computing device $D_{d}$ is denoted as $D_{d}=\{\varsigma_{d},\psi_{d}\}$, where $\varsigma_{d}$ is usually measured by device’s CPU [18]. The inherent heterogeneity of the 6G CPN is captured by the diverse characteristics of these three layers, where devices possess varying computing capacities, service prices, and are subject to different network conditions. Meanwhile, $\mathbb{T}=\{T_{1},\ldots,T_{t},\ldots,T_{T}\}$ denotes the set of all computing tasks. The tuple $T_{t}=\{\varphi_{t},\xi_{t},\tau_{t},d_{t}\}$ describes the computing task $T_{t}$. $\varphi_{t}$ and $\xi_{t}$ can be obtained using methods described in [19], such as graph analysis. Notations that will be used are presented in Table II.
TABLE II: Summary of Notations
@@
$\xi_{d}$ the average computing amount (in CPU cycles) of tasks assigned to $D_{d}$
@@
$\alpha_{d}$ the number of tasks processed per unit of $D_{d}$’s computing capacity